OpenStudy (anonymous):

"I am 3 times as old as Sue is," Frank said to Ann. "On the other hand, I am 15 years older than John while Sue is 1 year younger than John. How old am I?" If J represents John's age, then which of the following equations could be used to solve the problem? 3(J + 15) = J - 1 J + 15 = 3J - 1 J + 15 = 3(J - 1)

OpenStudy (anonymous):

It's the last one, because (J + 15) = Frank's age. So, three times Sue's age = Frank's age.

OpenStudy (anonymous):

Since he's 3 times as old as Sue, if we let Sue's age be S, then his age is 3S. Then we know that S=J-1, so his age is 3(J-1). We also know that he's 15 years older than John, so his age is also J+15. So we have J+15=3(J-1)